# FARIS: Fluid-Active-RIS

### v1

@@ Abstract
In this paper, we introduce a fluid-active reconfigurable intelligent surface (FARIS) that combines fluid-based port repositioning with per-element active amplification to enhance the performance of 6G network. To characterize the performance, we formulate an ergodic-rate maximization problem that jointly optimizes both the active amplification-reflection vector and the discrete selection of fluid active elements under practical hardware constraints. The problem is addressed via an alternating optimization (AO) framework, which progressively improves the rate. Complexity and convergence analyses that follow furnish deeper insight into the algorithmic operation and performance enhancement. Numerical results confirm that the proposed FARIS with AO framework consistently outperforms conventional FRIS/ARIS, delivering higher rates across diverse environments, often even when using fewer active elements or a smaller physical aperture.
@@ I Introduction
Recently, reconfigurable intelligent surfaces (RISs) have emerged as a key technology for sixth-generation (6G) wireless networks, primarily due to their capability to passively reconfigure the radio environment with minimal power consumption [1, 2]. An RIS typically comprises a large array of programmable, low-cost metasurface elements whose tunable reflection characteristics enable fine-grained manipulation of electromagnetic waves [3]. By coordinating the phase shifts across these elements, RISs can intentionally steer, focus, or scatter incident signals, thereby improving coverage, link reliability, spectral efficiency, and even localization performance across diverse wireless scenarios [4, 5]. Although RIS is recognized for its cost- and energy-efficiency in boosting wireless coverage and capacity, its practical deployment remains challenging. One critical issue is that since conventional RIS architectures rely solely on passive reflecting elements, it inevitably suffers from severe multiplicative fading, as their end-to-end path loss (PL) is the product of the Tx-RIS and RIS-Rx channel losses. This fundamental limitation significantly degrades performance in practical deployments [6, 7].
To address this fundamental physical bottleneck, recent research has introduced active-RIS (ARIS) as a new architecture for wireless communication systems [8, 9]. Unlike conventional passive-RIS (PRIS) that merely reflects incident signals, ARIS incorporates reflection-type amplifiers into their elements, enabling them to amplify the reflected signals. Although this requires additional power consumption, the resulting gain can effectively compensate for the severe double-reflection path loss, making ARIS a promising solution to the “multiplicative fading” problem. Several studies have further demonstrated the performance gains of ARIS in various wireless settings, including joint beamforming optimization with sub-connected architectures [10], signal-to-noise-ratio (SNR)-oriented analyses under identical power budget with PRIS [11, 12], and average sum-rate maximization [13] and power-minimizing [14] frameworks under partial channel-state-information (CSI) conditions, highlighting its potential to fundamentally enhance coverage, reliability, and spectral efficiency.
@@
In this context, the concept of fluid antenna system (FAS) has emerged as a compelling solution [24, 25]. Unlike conventional fixed-structure antennas, FAS offers shape- and position-reconfigurability enabled by flexible materials and architectures [26, 27], thereby introducing additional physical-layer DoFs that can be exploited for performance enhancement [28]. Since the FAS is first introduced [24], several studies have been made to understand the various performance limits of FAS [28, 29, 30], enable channel estimation schemes for FAS [31, 32, 33], and apply with several other applications including RIS [34, 35], integrated-sensing-and-communications (ISAC) [36, 37], and direction-of-arrival (DoA) estimations [38, 39]. Taken together, these studies establish FAS as a promising paradigm that surpasses conventional antenna systems by unlocking higher spatial diversity, and by offering a versatile platform that can be seamlessly integrated into diverse 6G wireless functionalities.
@@
This naturally raises the question of whether combining FRIS with the signal-amplification capability of ARIS could unlock even greater benefits. Importantly, however, our proposed architecture is not a mere system-level stacking of ARIS and FRIS mechanisms. Rather, it introduces a fundamentally new form of metasurface in which each active reflecting element functions as a fluidic unit, capable of providing controllable amplification while dynamically repositioning within its assigned subregion. In other words, this fluid-active behavior is intrinsically integrated into the surface itself.
In particular, we introduce fluid-active-RIS (FARIS), a novel paradigm that integrates the signal-amplification functionality of ARIS with the spatial reconfigurability of FRIS. Unlike FRIS and static ARIS, FARIS equips the surface with a fluidic-active-element that can simultaneously amplify and reflect incoming signals while freely adjusting its position within a predefined aperture. This dual capability allows the surface to adapt its spatial energy distribution in real time, thereby enabling (i) position-adaptive spatial-DoF extraction and (ii) compensation of multiplicative fading through active gain. As a result, FARIS achieves a fundamentally stronger form of environment control, offering a powerful and versatile platform for 6G wireless network.
The main contributions are summarized as follows:
@@
    We introduce the FARIS framework for a downlink setting and develop a unified formulation for joint port selection and amplification-phase design. By explicitly modeling the spatial correlation among activated FARIS ports and incorporating a practical ARIS reflection-power constraint, we cast an ergodic-rate maximization problem that simultaneously optimizes the amplification-reflection vector and the discrete configuration of fluid-reconfigurable ports.
@@
    To address the mixed-integer non-convex nature of the problem, we develop an alternating-optimization (AO) framework for design on FARIS architecture. With the port configuration fixed, we employ sample average approximation (SAA) together with a quadratic fractional transform to convert the objective into a sequence of convex conic subproblems in a lifted matrix variable. When the resulting solution is not rank-one, Gaussian randomization and magnitude projection are applied. We additionally establish that this inner-loop procedure guarantees monotonic improvement of the SAA objective and convergence to a stationary point.
@@
    For discrete port selection under a cardinality constraint on the activable FARIS elements, given the current amplification-reflection vector, we propose a cardinality-constrained cross-entropy method (CEM) that draws Bernoulli activation samples and updates an independent Bernoulli distribution to fit the statistics of the high-rate elite subset. We derive a closed-form CEM update, showing that the optimal activation probabilities are parametrized by a single Lagrange multiplier and that the resulting mapping admits a unique solution enforcing the cardinality budget. The smoothed update is proven to monotonically increase the empirical log-likelihood, and yields a non-decreasing sequence of ergodic rates.
@@
    Building on these, we integrate the active-reflection design and the port selection into a unified AO framework. We prove that each outer iteration produces a non-decreasing ergodic-rate sequence by leveraging the monotonicity of both the lifted conic subproblem and the likelihood-improving CEM update. Consequently, the overall AO procedure is guaranteed to converge to a finite limit, providing a principled mechanism for jointly optimizing continuous amplification-phase variables and discrete fluid-port configurations. We conduct a detailed complexity analysis of each algorithmic block, where the resulting expressions characterize how the overall computational cost scales, thereby confirming the polynomial-time implementability of the FARIS with proposed AO framework.
  5. 5.
    Extensive Monte-Carlo simulations are performed over a wide range of system parameters, including transmit power, total number of FARIS elements, and normalized aperture. The results show that the proposed FARIS architecture with AO framework exhibits fast convergence within a modest number of outer AO iterations and achieves near-optimal ergodic rate with only marginal loss compared to brute-force search (BFS), while drastically reducing computational burden. Moreover, FARIS consistently outperforms conventional FRIS and ARIS baselines, achieving significantly higher rates even with fewer active elements or smaller apertures.
To the best of our knowledge, this is the first work to propose a new paradigm that tightly integrates fluid port repositioning with ARIS, complemented by analytical and numerical complexity analyses and extensive simulations demonstrating its near-optimality and clear performance advantages over benchmarks.
@@ II System Model
where $j_{0}(\cdot)$ denotes the zero-order spherical Bessel function of the first kind, and $d_{ij}=d\sqrt{\left(\operatorname{mod}(i,M_{x})-\operatorname{mod}(j,M_{x})\right)^{2}+\left(\left\lfloor\frac{i}{M_{x}}\right\rfloor-\left\lfloor\frac{j}{M_{x}}\right\rfloor\right)^{2}}$ is the distance between the $i$ and $j$th elements. Following [44] and differing from [40], each FARIS element can be viewed as a port of a fluid antenna structure. Each element operates in one of two states: “on” and “off”. In the “on” state, the element actively interacts with incoming electromagnetic waves, modifying their amplitude and phase to realize the desired reflection. In the “off” state, the element is terminated by a matched load, effectively isolating it from the impinging wave and preventing any alteration of the signal.
@@
Out of the total $M$ candidate ports, only $M_{o}$ are activated. This selection process is represented as
@@
where $\mathbf{e}_{i}$ denotes the $i$th canonical basis vector (i.e., the $i$th column of $\mathbf{I}_{M}$). This construction implies that only the ports indexed by $\{i_{n}\}_{n=1}^{M_{o}}$ are active on the FARIS. To ensure unique activation, all columns of $\mathbf{S}_{M_{o}}^{\mathrm{T}}$ must be mutually distinct, preventing redundant port selection.
@@
and with the power budget $P_{\max}$, the hardware constraint is
$P_{\mathrm{RIS}}\leq P_{\max}.$ (9)
@@ III-A1 Update of $\mathbf{V}$ ($\leftrightarrow\mathbf{v}$) for fixed $\{y_{s}\}$
where $\{\mathbf{z}_{n}\sim\mathcal{CN}(\mathbf{0},\mathbf{I}_{M_{o}})\}$ are i.i.d. standard circularly symmetric complex Gaussian random vectors. For each ${\mathbf{v}}_{n}$, we form the rank-one matrix $\mathbf{V}_{n}\triangleq{\mathbf{v}}_{n}{\mathbf{v}}_{n}^{*}$, and solve (25) with $\mathbf{V}$ replaced by fixed $\mathbf{V}_{n}$, which is convex in $\{w_{s},\xi_{s}\}$. Let the optimum be
$\tilde{R}_{n}\triangleq\frac{1}{S}\sum_{s=1}^{S}\log_{2}\big(1+\xi_{s}^{[n]}\big),$ (27)
where $\xi_{s}^{[n]}$ is the optimizer with respect to $\xi_{s}$. We then select the best randomized sample index as
$n^{\star}=\operatorname*{\arg\max}_{n\in\{1,\cdots,N_{\mathrm{rand}}\}}\tilde{R}_{n}.$ (28)
We then enforce ${\mathbf{v}}_{n^{\star}}$ (or $\mathbf{v}_{0}$) by element-wise magnitude projection onto $0,g_{\max}]$ to ensure feasibility [46]:
$\forall v_{i}^{\prime}\leftarrow\min\{{g_{\max},|v_{i}^{\prime}|}\}e^{j\angle v_{i}^{\prime}},$ (29)
where $v_{i}^{\prime}$ is the $i$th component of corresponding ${\mathbf{v}}_{n^{\star}}$ or $\mathbf{v}_{0}$.
@@ III-A2 Update of $y_{s}$ for fixed $\mathbf{V}$
8:
Project the magnitude of each element of $\mathbf{v}^{[t+1]}$ onto $0,g_{\max}]$
9:  else
10:   $\mathbf{v}^{[t+1]}\leftarrow$ principal eigenvector of $\mathbf{V}^{[t+1]}$
11:  end if
12:  for $s=1,\cdots,S$ do
13:   Update $y_{s}^{[t+1]}$ via (30) using $\mathbf{V}^{[t+1]}$
14:  end for
15:  Evaluate $\bar{R}_{S}^{[t+1]}$ by (15)
16:  $t\leftarrow t+1$
17:until $|\bar{R}_{S}^{[t]}-\bar{R}_{S}^{[t-1]}|<\epsilon_{v}$
@@
19:Output: $\mathbf{v}^{\dagger}(\leftrightarrow\mathbf{V}^{\dagger})$
@@
In Algorithm 1, we initialize $\mathbf{v}^{[0]}(\leftrightarrow\mathbf{V}^{[0]})$ by drawing i.i.d. $\theta_{i}\sim\mathcal{U}[0,2\pi)$, set $\bar{\mathbf{v}}=g_{\max}[e^{j\theta_{1}},\cdots,e^{j\theta_{M_{o}}}]^{\mathrm{T}}$, and scale $\bar{\mathbf{v}}\rightarrow\mathbf{v}^{[0]}$ so that $\mathbf{v}^{[0]*}\mathbf{F}\mathbf{v}^{[0]}\leq P_{\max}$. This guarantees feasibility of (9) and $\forall|v_{i}^{[0]}|\leq g_{\max}$. Thereafter, we set $y_{s}^{[0]}=\frac{\sqrt{\mathrm{tr}(\mathbf{a}_{s}\mathbf{a}_{s}^{*}\mathbf{V}^{[0]})}}{\sigma_{0}^{2}+\mathrm{tr}(\mathbf{C}_{s}\mathbf{V}^{[0]})}~(\forall s)$.

### v2

# Fluid-Active-RIS (FARIS): A New Smart-Radio Paradigm with Joint Port Selection and Active Reflection Design

@@ Abstract
In this paper, we introduce a new wireless paradigm termed fluid-active reconfigurable intelligent surface (FARIS) that combines fluid-based port repositioning with per-element active amplification to enhance the performance of 6G networks. To realistically characterize the hardware operation, we first develop a circuit-level abstraction of the FARIS architecture and establish a practical power consumption model that captures both the logical control/switching power of candidate ports and the direct current (DC) bias power required for active reflection. Based on this model, we establish the FARIS signal model and formulate a corresponding ergodic-rate maximization problem that jointly optimizes the active amplification-reflection vector and the discrete selection of fluid-active elements under practical hardware constraints. The problem is addressed via an alternating optimization (AO) framework, which progressively improves the rate. Complexity and convergence analyses that follow furnish deeper insight into the algorithmic operation and performance enhancement. Numerical results confirm that the proposed FARIS with the AO framework consistently outperforms conventional baselines, delivering higher rates across diverse environments, often even when using fewer active elements or a smaller physical aperture.
@@ I Introduction
Recently, reconfigurable intelligent surfaces (RISs) have emerged as a key technology for sixth-generation (6G) wireless networks, primarily due to their capability to passively reconfigure the radio environment with minimal power consumption [41, 26]. An RIS typically comprises a large array of programmable, low-cost metasurface elements whose tunable reflection characteristics enable fine-grained manipulation of electromagnetic waves [14]. By coordinating the phase shifts across these elements, RISs can intentionally steer, focus, or scatter incident signals, thereby improving coverage, link reliability, spectral efficiency, and even localization performance across diverse wireless scenarios [18, 19, 16, 15]. Although RIS is recognized for its cost- and energy-efficiency in boosting wireless coverage and rate, its practical deployment remains challenging. One critical issue is that since conventional RIS architectures rely solely on passive reflecting elements, it inevitably suffers from severe multiplicative fading, as their end-to-end path loss (PL) is the product of the Tx-RIS and RIS-Rx channel losses. This fundamental limitation significantly degrades performance in practical deployments [4, 9].
To address this fundamental physical bottleneck, recent research has introduced active-RIS (ARIS) as a new architecture for wireless systems [1, 52]. Unlike conventional passive-RIS (PRIS) that merely reflects incident signals, ARIS incorporates reflection-type amplifiers into their elements, enabling them to amplify the reflected signals. Although this requires additional power consumption, the resulting gain can effectively compensate for the severe double-reflection path loss, making ARIS a promising solution to the “multiplicative fading” problem. Several studies have further demonstrated the performance gains of ARIS in various wireless settings, including joint beamforming optimization with sub-connected architectures [56], signal-to-noise-ratio (SNR)-oriented analyses under identical power budget with PRIS [53, 24], and average sum-rate maximization [47] and power-minimizing [54] frameworks under partial channel-state-information (CSI) conditions, highlighting its potential to fundamentally enhance coverage, reliability, and spectral efficiency.
@@
In this context, the concept of fluid antenna system (FAS) has emerged as a compelling solution [40, 29]. Unlike conventional fixed-structure antennas, FAS offers shape- and position-reconfigurability enabled by flexible materials and architectures [35, 51], thereby introducing additional physical-layer DoFs that can be exploited for performance enhancement [28]. Since the FAS is first introduced [40], several studies have been made to understand the various performance limits of FAS [28, 39, 27], enable channel estimation schemes for FAS [44, 30], and apply with several other applications including RIS [49, 11], integrated-sensing-and-communications (ISAC) [55, 13], and direction-of-arrival (DoA) estimations [46, 45]. Taken together, these studies establish FAS as a promising paradigm that surpasses conventional antenna systems by unlocking higher spatial diversity, and by offering a versatile platform that can be seamlessly integrated into diverse 6G wireless functionalities.
@@
From this observation, we conceive that the advantages of spatial mobility and active reflection on RIS are not merely additive but fundamentally complementary. On the one hand, ARIS can compensate for severe cascaded path loss through signal amplification [52]; however, its amplification occurs over a fixed surface geometry, limiting the ability to concentrate amplification on spatially favorable locations. On the other hand, FRIS can adaptively select advantageous spatial positions across a larger aperture [33], but passive reflection alone cannot sufficiently overcome the inherent multiplicative-fading penalty [4]. Therefore, the key opportunity lies in a deeper integration where each selected element simultaneously possesses spatial mobility and controllable amplification. In such a design, the surface can jointly optimize where the signal interaction occurs and how strongly the signal is reinforced, thereby creating a coupled spatial-amplitude design space that does not exist in conventional FRIS or static ARIS architectures. Importantly, such an architecture should not be interpreted as a simple system-level stacking of FRIS and ARIS functionalities. Rather, it requires a fundamentally new metasurface structure in which each reflecting element operates as a fluidic active unit, capable of dynamically repositioning within its designated subregion while simultaneously applying controllable amplification and phase adjustment. Through this intrinsic integration, spatial selection and signal reinforcement become jointly controllable design variables, enabling the surface to adapt both the spatial distribution and strength of reflected energy.
Motivated by this insight, we propose a fluid-active reconfigurable intelligent surface (FARIS), a new RIS paradigm that tightly integrates the spatial reconfigurability of FRIS with the signal amplification capability of ARIS. Unlike conventional FRIS or static ARIS architectures, FARIS equips each selected element with fluidic mobility together with active reflection functionality, allowing it to amplify and reflect incident signals while freely adjusting its position within the available aperture. This dual capability enables FARIS to simultaneously exploit position-adaptive spatial degrees of freedom and mitigate multiplicative fading through controllable active gain. As a result, FARIS provides a significantly stronger level of environment control compared with existing RIS architectures, offering a versatile platform for future 6G wireless networks. The main contributions are summarized as follows:
@@
    We introduce a newly proposed FARIS framework for a downlink setting and develop a unified formulation for joint port selection and amplification-phase design. By explicitly modeling the spatial correlation among selected FARIS ports and incorporating a practical ARIS reflection-power constraint, we cast an ergodic-rate maximization problem that simultaneously optimizes the amplification-reflection vector and the discrete configuration of fluid-reconfigurable ports. In addition, based on a circuit-level modeling of the FARIS architecture, we establish a practical power consumption model that captures both the control-circuit power and the direct current (DC) bias power of the active reflection modules, which enables a realistic evaluation of the overall power consumption.
@@
    To address the mixed-integer non-convex nature of the problem, we develop an alternating-optimization (AO) framework for design on FARIS architecture. With the port configuration fixed, we employ sample average approximation (SAA) together with a quadratic fractional transform to convert the objective into a sequence of convex conic subproblems in a lifted matrix variable. When the resulting solution is not rank-one, Gaussian randomization and magnitude projection are applied. We additionally establish that this inner-loop procedure guarantees monotonic improvement of the SAA objective and convergence to a stationary point. For discrete port selection under a cardinality constraint on the selectable FARIS elements, given the current amplification-reflection vector, we propose a cardinality-constrained cross-entropy method (CEM) that draws Bernoulli selection samples and updates an independent Bernoulli distribution to fit the statistics of the high-rate elite subset. We derive a closed-form CEM update, showing that the optimal selection probabilities are parametrized by a single Lagrange multiplier and that the resulting mapping admits a unique solution enforcing the cardinality budget. The smoothed update is proven to monotonically increase the empirical log-likelihood, and yields a non-decreasing sequence of ergodic rates.
@@
    Building on these, we integrate the active-reflection design and the port selection into a unified AO framework. We prove that each outer iteration produces a non-decreasing sequence with finite limit by leveraging the monotonicity of the subproblems in AO framework. We conduct a detailed complexity analysis of each algorithmic block, where the resulting expressions characterize how the overall computational cost scales, thereby confirming the polynomial-time implementability of the FARIS with proposed AO framework.
@@
    Extensive simulations are performed over a wide range of system parameters, including transmit power, total number of FARIS elements and candidates, and normalized aperture. The results show that the proposed FARIS architecture with the AO framework exhibits fast convergence within a modest number of outer AO iterations and achieves near-optimal ergodic rate with only marginal loss compared to brute-force search (BFS), while drastically reducing computational burden. Moreover, FARIS consistently outperforms conventional baselines, achieving significantly higher rates even with fewer active elements or smaller apertures. In addition, the power consumption analysis demonstrates that FARIS requires only a slightly higher power than ARIS due to the candidate port control, while providing substantially improved rate performance. Moreover, compared to the amplify-and-forward (AF) relay architecture, FARIS avoids the excessive hardware power consumption associated with complex architectures. As a result, FARIS achieves a more favorable rate-power trade-off among the considered architectures.
To the best of our knowledge, this is the first work to formalize FARIS as a new RIS paradigm that intrinsically couples fluid port reconfigurability with active reflection within a unified architecture. Supported by rigorous algorithmic design, detailed complexity characterization, and extensive simulations, our results show that FARIS is not merely a hybrid extension of existing RIS concepts, but a practically implementable and performance-enhancing paradigm that achieves near-optimal operation while consistently outperforming conventional FRIS and ARIS benchmarks.
@@ II System Model
where $j_{0}(\cdot)$ denotes the zero-order spherical Bessel function of the first kind, and $d_{ij}=d\sqrt{\left(\operatorname{mod}(i,M_{x})-\operatorname{mod}(j,M_{x})\right)^{2}+\left(\left\lfloor\frac{i}{M_{x}}\right\rfloor-\left\lfloor\frac{j}{M_{x}}\right\rfloor\right)^{2}}$ is the distance between the $i$ and $j$th elements. Following [42, 43] [^1] Although the cited model is for FRIS, we adopt it here since FARIS follows the same hardware architecture, with the difference being the additional amplification module for each candidate. and differing from [33], each FARIS element can be viewed as a port of a fluid antenna structure. Each element operates in one of two states: “on” and “off”; In the “on” state, the element is connected to the active reflection branch and actively interacts with the incident electromagnetic wave. Specifically, the impinging signal is processed through the controllable reflection circuitry, where its amplitude and phase can be adjusted according to the configured reflection coefficient. This enables the element to participate in the FARIS reflection process and contribute to shaping and amplifiying the composite reflected wave toward the desired direction. In contrast, in the “off” state, the element is terminated by a matched load through the switching network. In this case, the incident electromagnetic energy is absorbed rather than re-radiated, effectively isolating the element from the reflection process and preventing any modification of the impinging signal. Consequently, the element does not contribute to the reflected field and only incurs the baseline control-circuit power consumption. The schematic illustration of the FARIS architecture is given in Fig. 2, which serves as the basis for the power consumption model developed in the subsequent analysis.
Figure 2: Schematic illustration of the FARIS architecture with incorporating on-off control circuits and reflection-type amplifiers for the selected elements.
@@
Out of the total $M$ candidate ports, only $M_{o}$ are selected. This selection process is represented as
@@
where $\mathbf{e}_{i}$ denotes the $i$th canonical basis vector (i.e., the $i$th column of $\mathbf{I}_{M}$). This construction implies that only the ports indexed by $\{i_{n}\}_{n=1}^{M_{o}}$ are active on the FARIS. To ensure unique selection, all columns of $\mathbf{S}_{M_{o}}^{\mathrm{T}}$ must be mutually distinct, preventing redundant port selection.
@@
which accounts for both the amplified incident signal component and the amplified thermal noise emitted by the FARIS. In addition to $P_{\mathrm{RIS}}$, the FARIS operation also incurs the per-element hardware power consumption due to the selected fluid-active ports. Specifically, following the structure and circuit model in Fig. 2, the circuit power consumption of the FARIS is composed of two components: i) $P_{c}$: the logical control and switching power consumed by each candidate element, and ii) $P_{\mathrm{DC}}$: the DC bias power required for active reflection [24]. Since the logical control circuit exists for all $M$ candidates while the active reflection branch is given only for the selected $M_{o}$ FARIS ports, the total FARIS power consumption is modeled as
$P_{\mathrm{FARIS}}^{\mathrm{tot}}\triangleq MP_{c}+M_{o}P_{\mathrm{DC}}+\xi P_{\mathrm{RIS}},$ (9)
where $\xi\triangleq\frac{1}{\upsilon}$ with $\upsilon\in(0,1]$ denoting the amplifier efficiency. Therefore, under the total FARIS power budget $P_{\max,t}$, the feasible reflection design must satisfy $P_{\mathrm{FARIS}}^{\mathrm{tot}}\leq P_{\max,t}$, which is equivalent to:
$P_{\mathrm{RIS}}\leq\upsilon\big(P_{\max,t}-MP_{c}-M_{o}P_{\mathrm{DC}}\big)\triangleq P_{\max}.$ (10)
@@ III-A1 Update of $\mathbf{V}$ ($\leftrightarrow\mathbf{v}$) for fixed $\{y_{s}\}$
where $\{\mathbf{z}_{n}\sim\mathcal{CN}(\mathbf{0},\mathbf{I}_{M_{o}})\}$ are i.i.d. standard circularly symmetric complex Gaussian random vectors. For each ${\mathbf{v}}_{n}$, we first enforce feasibility with respect to the hardware constraints. Specifically, we apply an element-wise magnitude projection to satisfy $\{V_{ii}\leq g_{\max}^{2}\}$ in (26),
$\big[\mathbf{v}_{n}\big]_{i}\leftarrow\min\{g_{\max},|[\mathbf{v}_{n}]_{i}|\}e^{j\angle[\mathbf{v}_{n}]_{i}}~(i=1,\cdots,M_{o}),$ (28)
followed by a global power scaling
$\mathbf{v}_{n}\leftarrow\alpha_{n}\mathbf{v}_{n},~\alpha_{n}\triangleq\min\left\{1,\sqrt{\frac{P_{\max}}{\mathbf{v}_{n}^{*}\mathbf{F}\mathbf{v}_{n}}}\right\},$ (29)
which guarantees $\mathbf{v}_{n}^{*}\mathbf{F}\mathbf{v}_{n}\leq P_{\max}$ in (26). Let $\mathbf{v}_{n}^{\mathrm{feas}}$ denote the resulting feasible vector and define the corresponding rank-one matrix $\mathbf{V}_{n}\triangleq\mathbf{v}_{n}^{\mathrm{feas}}\mathbf{v}_{n}^{\mathrm{feas}*}$. For each $\mathbf{V}_{n}$, we evaluate the objective in (22), denoted by $\tilde{R}_{n}$, then select the best randomized candidate as
$n^{\star}=\operatorname*{\arg\max}_{n\in\{1,\cdots,N_{\mathrm{rand}}\}}\tilde{R}_{n},$ (30)
which corresponds to $\mathbf{v}_{n^{\star}}^{\mathrm{feas}}$.
@@ III-A2 Update of $y_{s}$ for fixed $\mathbf{V}$
8:  else
9:   $\mathbf{v}^{[t+1]}\leftarrow$ principal eigenvector of $\mathbf{V}^{[t+1]}$
10:  end if
11:  for $s=1,\cdots,S$ do
12:   Update $y_{s}^{[t+1]}$ via (31) using $\mathbf{V}^{[t+1]}$
13:  end for
14:  Evaluate $\bar{R}_{S}^{[t+1]}$ by (16)
15:  $t\leftarrow t+1$
16:until $|\bar{R}_{S}^{[t]}-\bar{R}_{S}^{[t-1]}|<\epsilon_{v}$
@@
18:Output: $\mathbf{v}^{\dagger}(\leftrightarrow\mathbf{V}^{\dagger})$
@@
In Algorithm 1, we initialize $\mathbf{v}^{[0]}(\leftrightarrow\mathbf{V}^{[0]})$ by drawing i.i.d. $\theta_{i}\sim\mathcal{U}[0,2\pi)$, set $\bar{\mathbf{v}}=g_{\max}[e^{j\theta_{1}}\cdots e^{j\theta_{M_{o}}}]^{\mathrm{T}}$, and scale $\bar{\mathbf{v}}\rightarrow\mathbf{v}^{[0]}$ so that $\mathbf{v}^{[0]*}\mathbf{F}\mathbf{v}^{[0]}\leq P_{\max}$. This guarantees feasibility of (10) and $\forall|v_{i}^{[0]}|\leq g_{\max}$. Thereafter, we set $y_{s}^{[0]}=\frac{\sqrt{\mathrm{tr}(\mathbf{a}_{s}\mathbf{a}_{s}^{*}\mathbf{V}^{[0]})}}{\sigma_{0}^{2}+\mathrm{tr}(\mathbf{C}_{s}\mathbf{V}^{[0]})}~(\forall s)$.